Accelerator Update for January 9, 2008 - January 11, 2008
Wednesday January 9
The midnight shift began with Operations monitoring the Tevatron (TeV) store 5835 (70.86E30) and the Antiproton Source stack (192.9E10), with no Recycler stash due to a Main Injector (MI) access for necessary repairs, with the Linac and Booster off and in access for repairs, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, and with the NuMI and SciBooNE / MiniBooNE (BNB) experiments also off due to MI and Booster accesses.
Operations prepared NuMI for an access at 8:08 AM.
RF experts completed their Booster work at 1:01 PM. All 19 RF stations were on.
Operations had Booster back on by 1:52 PM.
Operations reported at 1:56 PM that the MI vacuum leak was found on the upstream side of a RF cavity (MIRF8). Unfortunately, to weld the leak on the cavity they would have to pull the cavity into the isle, which put the proper repair of the leak beyond the timeframe of the access. Vacuum experts put a temporary “Vacseal” on the leak and they reported that the vacuum was holding.
At 2:34 PM, a Booster RF station (BRF12) tripped off. An RF expert investigated.
At 3:26 PM, Linac experts reported problems with the LRF2 switchtube replacement. The spare tube that was installed was found to have a vacuum problem and was being replaced.
An MI expert ended his collimator study around 4 PM.
A Linac expert reported at 4:44 PM that they were preparing to take beam.
At 4:55 PM, LRF2 tripped off for the first of two times in one minute. It reset.
MI RF experts reported at 5:09 PM that they were done with their RF work.
Operations established beam to Linac at 5:15 PM.
Operators began turning on MI at 5:21 PM.
A Linac expert began tuning the beam line at 5:24 PM.
At 6:07 PM, MI suffered a power supply interlock loop fault with volts to ground at MI-10 and MI-30. The fault reset.
Operations established beam to the Booster absorber at 6:34 PM.
Operations established beam to the MI absorber at 7:33 PM.
At 7:34 PM, Operations turned off all beam so that Booster experts could swap out a power supply regulator (B:SQL3).
Booster resumed sending beam at 8:01 PM.
Operations established beam to BNB at 8:10 PM.
A TeV expert began a Tevatron Electron Lens (TEL2) study at 8:17 PM.
Pbar resumed stacking at 8:18 PM.
At 8:22 PM, Operations reported a problem with the new Booster BEX power supplies. Booster experts reconfigured the system to run with the old supplies.
Operations established beam to NuMI at 9:15 PM.
At 9:33 PM, Operations turned off beam to NuMI and BNB so that Booster experts could investigate beam extraction problems.
Operations turned on the Pelletron at 9:59 PM.
The TEL2 studies ended at 10:21 PM. TeV experts began an end-of-store study.
At 11:54 PM, Operations reported a problem with a Pbar Debuncher RF station (DRF1). An expert said he’d come in and investigate.
Operations reported at midnight that the Recycler had stashed antiprotons once during the evening shift.
Thursday January 10
The midnight shift began with Operations monitoring store 5835 and the stash (172.6E10), with Pbar stacking (7.7E10), and with NuMI and BNB taking beam.
Operations terminated store 5835 at 12:04 AM.
At 12:09 AM, the BNB horn tripped off and would not reset. Operations called in an expert.
Operations prepared D0 for an access around 12:40 AM.
Pbar RF experts reported at 1:13 AM that they had made a temporary repair to DRF1-1 that should keep the tube running for the night, but the tube was drawing a lot of current and would have to be changed soon.
At 1:25 AM, LRF2 tripped off. It reset.
At 2:18 AM, MIRF8 tripped off and would not reset. An RF expert talked the operators through an extended reset.
At 3:11 AM, BRF14 tripped off. It reset.
Operations put the TeV through a dry squeeze at 4:13 AM.
Operations put the TeV into shot setup at 4:45 AM.
Operations reported at 5:26 AM that part of a stash, attempted at the beginning of this shift that didn’t make it into the Recycler, was injected into MI in the wrong spot. This explained why, when the MI extraction kickers fired, there was no beam sent out.
Operations reported that the BNB horn was fixed. The experiments resumed taking beam around 5:30 AM.
Operations established store 5837 at 6:58 AM with an initial luminosity of 184.56E30.
At 7:24 AM, the NuMI horn tripped off on a ground fault indication. It reset.
Operations reported at 7:59 AM that the Recycler had stashed four times during the midnight shift.
At 8:02 AM, KRF3 tripped off. It reset.
At 8:06 AM, MIRF2 tripped off. It reset.
At 9:54 AM, a Pbar Debuncher kicker (D:EKIK) module was not pulsing. Operations put Pbar in standby while an expert investigated.
At 9:59 AM, MIRF8 tripped off for the first of several times in a few minutes. An expert bypassed the station while he investigated; this allowed NuMI to continue taking beam.
At 10:09 AM, LRF2 tripped. It reset.
A TeV expert began a parasitic TEL2 study at 10:17 AM.
Pbar resumed stacking at 10:30 AM.
At 10:36 AM, MIRF2 tripped off. It reset.
At 10:37 AM, LRF2 tripped off. It reset.
The Neutron Therapy Facility (NTF) reported at 11:34 AM that they had treated their last patient for the day.
At 12:06 PM, MIRF2 tripped off and would not reset. Operators bypassed the station. Experts began replacing its power supply.
At 12:21 PM, the NuMI beam permit tripped off. It reset.
Pbar experts began a study at 12:22 PM.
The TeV TEL2 study ended at 1:09 PM.
At 1:19 PM, LRF2 tripped off. It reset.
At 3:28 PM, LRF2 tripped off. Experts investigated and then reset the station.
At 3:54 PM, KRF6 tripped off. It reset.
At 3:56 PM, LRF2 tripped off. It reset.
Operations reported at 7:59 PM that the Recycler had stashed three times during the day shift.
At 4:27 PM, Operations reported multiple Booster RF station tripped off. Everything reset.
At 4:30 PM, MI RF experts reported that MIRF2 also had a bad power amplifier and there were other problems that need further investigation, which the experts will check out on Friday.
At 9:54 PM, a Pbar sweeping magnet (M:USWSTA) PC failed. It was eventually reset.
At 10:25 PM, DRF1-6 tripped off. The station would not reset and was bypassed.
Operations reported at midnight that the Recycler had stashed six times during the evening shift.
Friday January 11
The midnight shift began with Operations monitoring store 5837 (43.93E30) and the stash (304.4E10), with Pbar stacking (28.0E10), and with NuMI and BNB taking beam.
Operations terminated store 5837 at 1:04 AM.
Operations put the TeV into shot setup at 1:23 AM.
At 1:47 AM, the Booster West Anode Power Supply tripped off. It reset.
At 1:52 AM, BRF9 tripped off. It reset.
Operations established store 5838 at 3:26 AM with an initial luminosity of 217.89E30.
At 6:16 AM, LRF2 tripped off. It reset.
Operations reported at 7:55 AM that the Recycler had stashed four times during the midnight shift.
The Plan for Friday and the Weekend
The plan for today and the weekend is to stack, stash, and store.
Comments and Suggestions
|last modified 01/14/2008 email Fermilab|
|Security, Privacy, Legal|